# Hyperon spin correlation in high-energy heavy-ion collisions

### v2

@@ Introduction.
which denotes the difference between the “same-sign” pairs ($\Lambda\Lambda$ or $\bar{\Lambda}\bar{\Lambda}$) and the “opposite-sign” pairs ($\Lambda\bar{\Lambda}$). We show in Figs. 3 and 4 its dependence on the azimuthal-angle difference and the rapidity difference between the two hyperons. When the SFF is taken into account, the $xx$, yyyy, and $zz$ components of $c_{\text{net}}$ peak at $\delta\phi=0$ or $\delta Y=0$ and approach zero when $\delta\phi$ or $\delta Y$ increase. In the absence of SFF, $c_{\text{net}}$ becomes 2-3 orders of magnitude smaller than in the case with SFF, shown in the insets in Figs. 3 and 4. This behavior arises because the SFF has an opposite effect on the polarizations of $\Lambda$ and $\bar{\Lambda}$, as described by Eq. (5), leading to opposite contributions to the spin correlations of $\Lambda\Lambda$ and $\Lambda\bar{\Lambda}$. In contrast, the polarizations induced by the thermal vorticity and the thermal shear tensor are identical for $\Lambda$ and $\bar{\Lambda}$, resulting in nearly the same values of $c_{\Lambda\Lambda}^{yy}$, $c_{\bar{\Lambda}\bar{\Lambda}}^{yy}$, and $c_{\Lambda\bar{\Lambda}}^{yy}$ when the SFF is switched off. The spin Hall effect could also result in a nonzero $c_{\text{net}}$, but its magnitude is much smaller than the contribution of SFF. Therefore $c_{\text{net}}$ is shown to be sensitive to the SFF and is almost independent from the other hydrodynamic fields.
@@
Figure 3: Components of the net spin correlation $c^{ab}_{\text{net}}$ as functions of the azimuthal-angle difference. Results with the SFF are shown by the orange, red, and cyan lines, while the other lines are results in absence of the SFF.
@@
Figure 4: Components of the net spin correlation $c^{ab}_{\text{net}}$ as functions of the rapidity difference. Notations are the same as in Fig. 2.
@@
The dependence of the net spin correlation on $\delta\phi$ or $\delta Y$ is a consequence of the position-momentum correlation of hyperons. Due to the transverse expansion of the QGP, the mean transverse position of hyperons emitted from the freeze-out hypersurface is approximately aligned with the transverse momentum, $\left\langle{\bf x}_{T}\right\rangle\propto\hat{\bf p}_{T}$. Similarly, the longitudinal expansion leads to a nearly linear dependence of $\left\langle z\right\rangle$ on the rapidity. Given such significant position-momentum correlations, a larger separation of two hyperons in momentum space corresponds to a larger distance between their production positions, which leads to weaker field-field and spin-spin correlations according to Eqs. (8) and (9). The $\delta\phi$ or $\delta Y$ dependence is similar to the $\Lambda\bar{\Lambda}$ spin correlation observed in p+p collisions [43], but arises from a different mechanism. In p+p collisions (or for high-momentum hyperons in A+A collisions [59, 60]), it is dominated by the parton fragmentation, while the correlation in this study is dominated by interactions between constituent quarks and the fluctuating SFF.
@@
Given the SFF fluctuation constrained by the experimental data on the $\phi$ meson spin alignment, we have provided quantitative predictions on the relative momentum dependence ($\delta\phi$ and $\delta Y$) of the net spin correlations of hyperons, which all decrease with their separation in momentum space. We also predicted the dependence of different spin correlation Components on the collision energy, which shows a pattern similar to the $\phi$ meson spin alignment. Provided that the typical space-time scale of the SFF fluctuation is comparable to the size of a light meson, $\sigma_{t}=\sigma_{x}=0.6$ fm, the net spin correlation is $\mathcal{O}(10^{-4})$. Larger values of $\sigma_{t}$ and $\sigma_{x}$ result in stronger spin correlations. For example, if $\sigma_{t}=\sigma_{x}=1.2$ fm, the correlation induced by SFF will become twice as large.
These predictions provide a perfect opportunity to test the mechanism of spin polarization in the QGP and constraints on the scale of the SFF fluctuation in high-energy heavy-ion collisions. For simplicity, we have neglected the internal motion of $s$ quark inside the hyperon in the constituent-quark coalescence model. Additionally, the classical electromagnetic fields are not considered, as they should be weak at the freeze-out stage of heavy-ion collisions. These effects, along with further refinements, can be addressed in future studies.
@@ Acknowledgments.
The authors thank Diyu Shen and Aihong Tang for enlightening discussions. This work is supported in part by the Italian Ministry of University and Research, project PRIN2022“Advanced probes of the Quark Gluon Plasma”, Next Generation EU, Mission 4 Component 1, X.-Y.W. is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048 and SAPIN-2024-00026] and in part by US National Science Foundation (NSF) under grant number OAC2004571. D.H.R. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through CRC-TR 211 “Strong-interaction matter under extreme conditions” - project number 315477589 - TRR 211. X.-N.W. is supported by NSFC under Grant No. 1193507, the Guangdong MPBAR with No. 2020B0301030008 and by the Alexander von Humboldt Foundation through a Humboldt Research Award.

### v3

@@ Introduction.
which denotes the difference between the “same-sign” pairs ($\Lambda\Lambda$ or $\bar{\Lambda}\bar{\Lambda}$) and the “opposite-sign” pairs ($\Lambda\bar{\Lambda}$). We show in Figs. 3 and 4 its dependence on the azimuthal-angle difference and the rapidity difference between the two hyperons. When the SFF is taken into account, the $xx$, yyyy, and $zz$ components of $c_{\text{net}}$ peak at $\delta\phi=0$ or $\delta Y=0$ and approach zero when $\delta\phi$ or $\delta Y$ increase. In the absence of SFF, $c_{\text{net}}$ becomes 1-2 orders of magnitude smaller than in the case with SFF, shown in the insets in Figs. 3 and 4. This behavior arises because the SFF has an opposite effect on the polarizations of $\Lambda$ and $\bar{\Lambda}$, as described by Eq. (5), leading to opposite contributions to the spin correlations of $\Lambda\Lambda$ and $\Lambda\bar{\Lambda}$. In contrast, the polarizations induced by the thermal vorticity and the thermal shear tensor are identical for $\Lambda$ and $\bar{\Lambda}$, resulting in nearly the same values of $c_{\Lambda\Lambda}^{yy}$, $c_{\bar{\Lambda}\bar{\Lambda}}^{yy}$, and $c_{\Lambda\bar{\Lambda}}^{yy}$ when the SFF is switched off. The spin Hall effect could also result in a nonzero $c_{\text{net}}$, but its magnitude is much smaller than the contribution of SFF. Therefore $c_{\text{net}}$ is shown to be sensitive to the SFF and is almost independent from the other hydrodynamic fields.
@@
Figure 3: Components of the net spin correlation $c^{ab}_{\text{net}}$ as functions of the azimuthal-angle difference in the range $|Y|<1$, $p_{T}\in(0.5,\,3)$ GeV/c for Au+Au collisions at $\sqrt{s_{\text{NN}}}=27$ GeV. Results with the SFF are shown by the orange, red, and cyan lines, while the other lines are results in absence of the SFF.
@@
Figure 4: Components of the net spin correlation $c^{ab}_{\text{net}}$ as functions of the rapidity difference. Computational conditions and notations are the same as in Fig. 3.
@@
The dependence of the net spin correlation on $\delta\phi$ or $\delta Y$ is a consequence of the position-momentum correlation of hyperons. Due to the transverse expansion of the QGP, the mean transverse position of hyperons emitted from the freeze-out hypersurface is approximately aligned with the transverse momentum, $\left\langle{\bf x}_{T}\right\rangle\propto\hat{\bf p}_{T}$. Similarly, the longitudinal expansion leads to a nearly linear dependence of $\left\langle z\right\rangle$ on the rapidity. Given such significant position-momentum correlations, a larger separation of two hyperons in momentum space corresponds to a larger distance between their production positions, which leads to weaker field-field and spin-spin correlations according to Eqs. (8) and (9). The $\delta\phi$ or $\delta Y$ dependence is similar to the $\Lambda\bar{\Lambda}$ spin correlation observed in p+p collisions [STAR:2025njp], but arises from a different mechanism. In p+p collisions (or for high-momentum hyperons in A+A collisions [Li:2023qgj, Huang:2024awn]), it is dominated by the parton fragmentation, while the correlation in this study is dominated by interactions between constituent quarks and the fluctuating SFF.
@@
Given the SFF fluctuation constrained by the experimental data on the $\phi$ meson spin alignment, we have provided quantitative predictions on the relative momentum dependence ($\delta\phi$ and $\delta Y$) of the net spin correlations of hyperons, which all decrease with their separation in momentum space. We also predicted the dependence of different spin-correlation components on the collision energy, which shows a pattern similar to the $\phi$ meson spin alignment. Provided that the typical space-time scale of the SFF fluctuation is comparable to the size of a light meson, $\sigma_{t,x}=0.6$ fm, the net spin correlation is $\mathcal{O}(10^{-4})$. Larger values of $\sigma_{t}$ and $\sigma_{x}$ result in stronger spin correlations. A simple estimate for the required statistical precision shows that at least $10^{8}$ pairs are needed for measuring such a tiny effect, which can be achieved in the BES-II program at RHIC.
These predictions provide a perfect opportunity to test the mechanism of spin polarization in the QGP and constraints on the scale of the SFF fluctuation in high-energy heavy-ion collisions. It will provide new information about the short-range correlation in the QGP and the non-perturbative behaviour of strongly interacting matter at high temperature and density. For simplicity, we have neglected the internal motion of the $s$ quark inside the hyperon in the constituent-quark coalescence model. Additionally, the classical electromagnetic fields are not considered, as they should be weak at the freeze-out stage of heavy-ion collisions. These effects, along with further refinements, can be addressed in future studies.
@@ Acknowledgments.
The authors thank Diyu Shen and Aihong Tang for enlightening discussions. This work is supported in part by the Italian Ministry of University and Research, project PRIN2022“Advanced probes of the Quark Gluon Plasma”, Next Generation EU, Mission 4 Component 1, X.-Y.W. is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048 and SAPIN-2024-00026] and in part by US National Science Foundation (NSF) under grant number OAC2004571. D.H.R. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through CRC-TR 211 “Strong-interaction matter under extreme conditions” - project number 315477589 - TRR 211. X.-N.W. is supported by NSFC under Grant No. 1193507, the Guangdong MPBAR with No. 2020B0301030008 and by the Alexander von Humboldt Foundation through a Humboldt Research Award. The data that support the findings of this article are openly available [sheng_2025_17938737].
Supplemental Material for
“Hyperon spin correlation in high-energy heavy-ion collisions”
S1 Measuring spin correlation in experiments
In experiments, the spin polarization of the $\Lambda$ hyperon is extracted from its weak decay channel $\Lambda\rightarrow p+\pi^{-}$. The angular distribution of the daughter proton follows
$\frac{dN}{d\Omega^{\ast}}=\frac{1}{4\pi}\left(1+\alpha_{\Lambda}\frac{\boldsymbol{\mathcal{P}}^{\text{rest}}_{\Lambda}\cdot{\bf p}^{\ast}}{|{\bf p}^{\ast}|}\right)\,$, (S1)
where $\boldsymbol{\mathcal{P}}^{\text{rest}}_{\Lambda}$ is the spin-polarization vector of $\Lambda$, $\bf p$$^{\ast}$ is the polar angle of the proton momentum in the rest frame of the $\Lambda$, and $\alpha_{\Lambda}$ denotes the decay constant. This relation allows one to measure the polarization through
$\boldsymbol{\mathcal{P}}^{\text{rest}}_{\Lambda}=\frac{3}{\alpha_{\Lambda}}\left\langle\hat{\bf p}^{\ast}\right\rangle\,.$ (S2)
where $\hat{\bf p}^{\ast}\equiv{\bf p}^{\ast}/|{\bf p}^{\ast}|$ is the unit vector along the direction of $\bf p$$^{\ast}$. In analogy to this approach, the hyperon spin correlation introduced in Eq. (2) is measured from the decay products via
$c^{ab}_{12}=\frac{9}{\alpha_{1}\alpha_{2}}\left\langle\cos\theta_{1,a}\cos\theta_{2,b}\right\rangle\,$, (S3)
where $1,2$ refer to $\Lambda$ or $\bar{\Lambda}$, and $a,b=x,y,z$ denote three spatial directions. Here, $\cos\theta_{1,a}\equiv\hat{\bf p}_{1}^{\ast}\cdot{\bf e}_{a}$ and $\cos\theta_{2,b}\equiv\hat{\bf p}_{2}^{\ast}\cdot{\bf e}_{b}$ with $\hat{\bf p}^{\ast}_{i}$ being the unit vector of the daughter-proton’s momentum from the decay of the $i$-th hyperon, and ${\bf e}_{x,y,z}$ the unit vectors along the $x,y,z$ axes. The parameters $\alpha_{1}$ and $\alpha_{2}$ are the decay constants for hyperons 1 and 2, respectively.
S2 Spin polarization of quark/antiquark
In the quark-gluon plasma described by a relativistic fluid, quarks and antiquarks can be polarized by various sources. In this work, we take into account the contributions from the thermal vorticity, the thermal shear tensor, the spin Hall effect, and the SFF. The average polarization as a function of phase-space coordinates is defined as
$P^{\mu}(x,p)\equiv n^{\mu}(x,p)\frac{f_{+}(x,p)-f_{-}(x,p)}{f_{+}(x,p)+f_{-}(x,p)}\;,$ (S4)
where $n^{\mu}(x,p)$ denotes the unit vector along the polarization direction for a particle (or antiparticle) with momentum $p$ at position $x$, and $f_{\pm}(x,p)$ are the number densities of the particle (or antiparticle) with spin projection $\pm 1/2$ along $n^{\mu}$. As shown in Eq. (4), the average polarization can be decomposed into several contributions, where
$P_{\omega}^{\mu}(x,p)$ $=-\frac{1}{4m}\epsilon^{\mu\nu\rho\sigma}p_{\nu}\varpi_{\rho\sigma}\;,$
$P_{\text{shear}}^{\mu}(x,p)$ $=-\frac{1}{2m(u\cdot p)}\epsilon^{\mu\nu\rho\sigma}p_{\nu}u_{\rho}\xi_{\sigma\lambda}p^{\lambda}\;,$
$P_{\text{SHE}}^{\mu}(x,p)$ $=\pm\frac{1}{2m(u\cdot p)}\epsilon^{\mu\nu\rho\sigma}p_{\nu}u_{\rho}\partial_{\sigma}\frac{\mu_{s}}{T}\;,$ (S5)
with $P_{\phi}^{\mu}$ given in Eq. (5). Here, $p^{\mu}=(\sqrt{{\bf p}^{2}+m^{2}},{\bf p})$ is the four-momentum for the considered particle (or antiparticle), while $u\cdot p$ is its energy in the frame comoving with the fluid. The explicit expression for $P_{\omega}^{\mu}(x,p)$ can be found in the review [Becattini:2024uha], where $S^{\mu}$ in this reference corresponds to $P^{\mu}/2$ in our notation. The shear-induced polarization has been formulated differently in different papers [Becattini:2021suc, Becattini:2021iol, Liu:2021uhn, Fu:2021pok]. In this work, we take the frame vector in $P_{\text{shear}}^{\mu}$ to be $u^{\mu}$, which is consistent with the formula in Ref. [Fu:2021pok].
S3 Estimate of the correlation length
When calculating the spin correlation induced by SFFs, we have employed a Gaussian smearing function (10) to suppress long-distance field correlations. This introduces $\sigma_{t}$ and $\sigma_{x}$ as the typical temporal and spatial scales, which can also be interpreted as the typical correlation length of the SFF (the $\phi$ field in this work). From the theoretical side, there is no precise model input for $\sigma_{t,x}$, so here we provide several estimates for the correlation length. The SFF interaction provides an effective description of the strong interaction between particles at the freeze-out stage, which lies intermediate between the hadronic interaction at low temperature and the gluonic interaction at high temperature. It is therefore natural to expect that $\sigma_{t,x}$ falls between the correlation length in a hadron gas and that of pure QCD. Several possible estimates are:
  • 1.
    In QCD at low energies, the hadronic interactions are predominantly mediated by mesons. Since mesons have a nonzero spatial extent rather than being point-like particles, it is reasonable to assume that the SFFs also extend over the meson’s spatial extent. The corresponding correlation length can thus be estimated by the charge radii of mesons: 0.66 fm for pions [ParticleDataGroup:2024cfk], 0.56 fm for kaons [ParticleDataGroup:2024cfk], and 0.61 for $\rho$ mesons [Hawes:1998bz].
  • 2.
    Alternatively, one can estimate the correlation length using the typical interaction scale, i.e., the Compton wavelength $\lambda\sim 1/m$. Taking $m$ as the vacuum masses of pion, kaon, or $\phi$ meson, the related scales are 1.4 fm, 0.4 fm, and 0.2 fm, respectively.
  • 3.
    In hot QCD matter where the strong interaction is mediated by gluon fields, the correlation length of color electric fields can be estimated via the Debye screening length [Blaizot:2001nr], $\lambda_{D}=1/m_{D}\sim 1/(gT)$. Taking the freeze-out temperature $T\approx 160$ MeV and the strong coupling constant $g\approx 2$, the Debye screening length is $\lambda_{D}\approx 0.62$ fm.
  • 4.
    In the CGC/Glasma framework, the transverse correlation length of the color fields is parametrized by the inverse saturation scale, $\lambda\sim 1/Q_{s}$ [Gelis:2010nm]. The typical value $Q_{s}\sim$ 1 GeV corresponds to $\lambda\sim$ 0.20 fm.
  • 5.
    According to lattice-QCD calculations, the correlation length of the gluon field is around 0.2 – 0.3 fm, or even shorter [campostrini1986g, DIGIACOMO2002319].
In our work, the spin correlation is calculated at the freeze-out stage of the collision, which lies intermediate between a hadron gas and a pure QCD matter. Therefore, the correlation length of SFF used here is expected to be neither as short as that of gluon field nor as long as the pion Compton wavelength. Based on the above considerations, we adopt a conservative estimate for the correlation length of $\sigma_{t,x}=0.6$ fm, with an uncertainty of $0.3$ fm.
S4 $\delta p_{T}$-dependence of spin correlation
Figure S1: The net spin correlation as a function of $\delta p_{T}\equiv|p_{T,1}-p_{T,2}|$. Computational conditions are the same as in Fig. 3.
We show the net spin correlation as a function of the transverse-momentum difference in Fig. S1, where $\delta p_{T}\equiv|p_{T,1}-p_{T,2}|$. In the presence of SFFs, the $xx$ and $yy$ components of the net spin correlation increase with $\delta p_{T}$, while the $zz$ component decreases. Contributions from hydrodynamic fields are found to be two orders of magnitude smaller than those from the SFFs, as illustrated in the inset of Fig. S1.
S5 Effect of spin Hall effect to spin correlation
The spin Hall effect, originating form the gradient of $\mu_{s}/T$, has opposite contributions to the polarization of quark and antiquark, as shown in Eq. (S2). Consequently, it could induce $c_{\text{net}}$ in a similar way as the SFFs. We show in Figs. S2 and S3 its contributions to $c_{\text{net}}$ as functions of $\delta\phi$ and $\delta Y$, respectively. The resulting correlations are of order $\mathcal{O}(10^{-6})$, which is significantly smaller than those induced by the SFFs.
Figure S2: The contribution of SHE to the net spin correlation as a function of $\delta\phi$. Computational conditions are the same as in Fig. 3.
Figure S3: The contribution of SHE to the net spin correlation as a function of $\delta Y$. Computational conditions are the same as in Fig. 3.